# Statistics within UV–Visible Absorption Spectrum of Ethanolic Azobenzene

### v1

@@ Abstract
We report a statistical simulation of the UV–visible absorption spectrum of trans-azobenzene in ethanol solution. Due to intermolecular interactions, the used explicit solvent environment necessitates accounting for numerous transitions for a spectrum covering the two energetically lowest lines, S1 and S2. Furthermore, the spectrum manifests vast variation as a function of the underlying local structure, in conjunction with previous observations for spectra of liquids in the X-ray regime. We disentangle the complex structure–spectrum relationship using a machine learning-based method known as the emulator-based component analysis. This structural decomposition outperforms commonly used principal component analysis in explained spectral variation and reveals a small subset of latent structural variables responsible for the total spectral variance. Among other structural characteristics, blueshifting of the S2 line occurs with fewer hydrogen bonds with the ethanol solvent, and a contracted N=N bond within the C–N=N–C bridge. The observed structural dependence of the absorption spectrum thus implies an overrepresentation of certain structural classes after a photoexcitation, potentially significant for the subsequent nuclear dynamics, photophysics, and photochemistry.
A liquid system exhibits a wealth of local atomistic structures, which is reflected as strong variation in their X-ray [1, 2, 3, 4, 5] and UV–visible [6, 7, 8] spectra. In general, sampling of the ensemble is required, as the effect is not necessarily captured by just a few representative structures and their respective spectra [9]. On one hand, statistical variation of the spectra implies pre-selection of structures for, e.g., absorption at a given wavelength. On the other hand, this dependence allows for interpreting the details of spectra for the purposes of characterization if the spectrally relevant structural degrees of freedom can be identified. In a liquid system, however, this is not completely straightforward. Fortunately, extensive simulations enable an in-depth study of the phenomenon when combined with analysis techniques capable of transforming numerical results into human-interpretable forms [10, 11, 12, 13].
@@
We investigate statistical variation in the UV–visible spectrum of trans-azobenzene in ethanol solution using machine learning-based analysis of data based on sampling by molecular dynamics (MD). Photoisomerization between the trans and the cis forms [14] of the molecule is an archetypal topic of photochemistry and a mechanism utilized in applications [15, 16]. We sampled 30 000 structures from extended tight binding MD (xTBMD) using the GFN1-xTB method [17, 18] implemented in the CP2K software [19] v. 2024.2. In the simulations, a trans-azobenzene molecule was dissolved in 159 ethanol molecules in a cubic box of $L$ = 25.04 Å for the density 0.7840 g/cm3 interpolated for pure ethanol liquid at 300 K [20]. After equilibration of 20 ps, we sampled the canonical ensemble at 300 K with a 10 fs interval from a production run of 300 ps. We used a global Nosé–Hoover thermostat [21, 22] with chain length 3, a time constant $\tau$ of 1000 fs, and an MD timestep of 0.5 fs. For evaluation of spectra consisting of 100 transitions on each of the 30 000 structures, we applied time-dependent density functional theory (TD-DFT) implemented in ORCA v.6.0.0 [23]. We explicitly included ethanol within the cutoff radius of 3.0 Å, and simulated this cluster in a conductor-like polarizable continuum model [24]. We used the PBE exchange-correlation potential [25], the def2-SVP basis set [26] and def2/J auxiliary basis set [27] in the spectral simulations. We evaluated the cross section $\sigma_{i}$ of each transition $i$ for a randomly oriented sample using the transition matrix element in the velocity form, and then convolved this stick spectrum to differential cross section $\mathrm{d}\sigma/\mathrm{d}E$ in terms of energy $E$ by a Gaussian of 0.2 eV full width at half maximum. Finally, we transformed the differential cross section into wavelength $\lambda$ scale for $\mathrm{d}\sigma/{\mathrm{d}\lambda}$. We use a constant width in energy, because a constant width in wavelength makes the lines artificially broad for short wavelengths.
Figure 1: Results of the spectrum calculations and variance captured by different dimensionality reduction techniques. a: Simulated ensemble average UV–visible absorption spectrum for trans-azobenzene in ethanol, together with an experiment reproduced from Ref. 28. The blue-shaded area depicts the standard deviation of the data set, indicating significant statistical variation. The gray-shaded areas depict the two regions of interest (ROIs I and II) defined for the subsequent analysis. Computational results have been shifted $-50.6$ nm to match the experiment. Additionally, the figure presents an example local structure prepared with Jmol [29]. b: Spectral R2 score as a function of the rank of emulator-based component analysis (ECA) and principal component analysis (PCA), evaluated using the neural network emulator after the respective dimensionality reduction. The first two ECA components explain a large portion of the spectral ROI variance, while the higher components show rapidly diminishing R2 score gains. Structural decomposition by PCA covers significantly less spectral variance for the same rank. c: Structural R2 score as a function of decomposition rank. The ECA vectors cover a small fraction of the structural variance compared to structural PCA. The result indicates that explaining structural variation alone does not explain the spectral variation.
Figure 1a shows the simulated ensemble-average spectrum shifted by $-50.6$ nm in comparison to the experimental one by Nägele and coworkers [28]. We apply the molecular notation S1 and S2 for the characteristic features in this spectrum, but note that in our simulation for the gas phase in the optimized geometry, the S2 line already consists of three transitions. In liquid, we find the photoexcited states of trans-azobenzene to be mixed with states of the ethanol, which results in a wealth of possible final states to be excited – and to be simulated with explicit solvation. In our data, the lines S1 and S2 typically consist of several notably intense transitions and numerous weaker ones. The S1 feature (390 nm–520 nm) is almost unobservable, and we focus on the S2 peak (270 nm–390 nm). We observe significant statistical variation between the spectra of individual local structures, as shown by the standard deviation $\sigma$ in Figure 1a. For the following analysis, we define two regions of interest (ROI), split at the maximum of the S2 peak. Then, we study the mean differential cross section within each ROI as a function of the structure. We first train a neural network to predict the two z-score standardized ROI values for a structure, and then utilize this emulator in emulator-based component analysis (ECA) [10] for dimensionality reduction to identify the decisive structural characteristics behind the observed spectral variation.
@@
We pre-process the structural information using a local version of the many-body tensor representation (LMBTR) [30] to obtain a vector $\mathbf{d}$ for each of the data points. We then apply z-score standardization for $\mathbf{\tilde{d}}$, the vector input to a neural network. The LMBTR descriptor consists of Gaussian-smeared element-wise radial distance distributions for multiple centers in the system. As these centers, we use all the nitrogen and carbon sites in the azobenzene molecule, and eight additional virtual centers located perpendicular to a plane defined by nearby atoms in the azobenzene molecule. The first four virtual centers lie ±$\pm 2$ from each nitrogen atom. The latter four centers lie ±$\pm 2$ perpendicular from the arithmetic mean position of the para and ortho carbon atoms of each ring. We ran a joint search [12] of the LMBTR and neural network hyperparameters, choosing the best-performing combination (for details, see SI). The final model had an R2 score (coefficient of determination) of 0.891 on a test set of 5 000 points, split from the data pool of 30 000 points prior to model selection, and thus previously unseen by the model. This performance is similar to that observed for X-ray spectra [5, 12, 13]. In addition to good ML and ECA performance, demonstrated for X-ray spectra, the LMBTR descriptor is also interpretable by a human.
@@
As seen in Figure 1b, rapid convergence towards the emulator performance limit is observed when emulation is carried out on data points projected onto the low-dimensional subspace. The ECA method is analogous to principal component analysis (PCA) with the key difference being that the optimization target of the former is the explained ROI intensity variance. In contrast, applying PCA to the structural data points $\mathbf{\tilde{d}}$ aims at finding the decisive components of structural variance, not necessarily reflected by the spectral response as seen in Figure 1b. Vectors from structural PCA cover significantly less spectral variance after the corresponding evaluation. Perhaps even more notably, the dominant structural degree of freedom is spectrally irrelevant. The different optimization goal of the methods is also evident in the structural R2 score as a function of the decomposition rank, presented in Figure 1c, which shows PCA to surpass the ECA decomposition. Thus, the majority of the spectral ROI intensity variance is explained by a minority of structural variance. We note that another method, using partial least squares fitting, has previously been compared to ECA, with the latter excelling in this comparison [10].
@@
Next, we turn to the analysis of the first two ECA component vectors, that indicate the structural changes responsible for most spectral ROI intensity variation. We first analyze the spectral effect, and then the associated structural changes. We define two spectral properties: (i) ROI value difference (ROI I – ROI II), and (ii) total ROI value sum. These properties can be directly linked in the two latent coordinates $t_{1}$ and $t_{2}$, as indicated in Figure 2. The spectrally most significant structural degree of freedom $t_{1}$ corresponds to the ROI value difference (Figure 2a), whereas the second most significant degree of freedom $t_{2}$ captures the ROI value sum (Figure 2d). The two properties are practically independent of each other, illustrated by the low cross-correlation in Figures 2b and 2c. From Figures 2a and 2d we conclude that ROI I value is high when $t_{1}$ is high and $t_{2}$ is low. Similarly, ROI II is high when both $t_{1}$ and $t_{2}$ are low. Thus, the structural changes causing a blueshift of the S2 line are indicated by the latent coordinate $t_{1}$, as the effects from the second ECA component cancel out. For a structural interpretation of such effect, we calculate the difference 10 $\mathbf{v}_{1}$ for a well-expected change in this coordinate, and then transform the respective ECA component vector to the original descriptor space. A further division by the squared distance to a corresponding center yields site-wise and element-wise curves with radial behavior proportional to that of the radial distribution function.
@@
Figure 2: The relationship between spectral properties and the latent coordinates from ECA using the test set with the corresponding Pearson’s correlation coefficients $r$. a: The difference between the two ROI values is highly correlated with the spectrally most significant structural variable $t_{1}$. b: The correlation between ROI value difference and the second most significant structural variable $t_{2}$ is weak. c: The total ROI value sum is weakly correlated with $t_{1}$, but d: highly correlated with $t_{2}$. For details, see text.
@@
Figure 3 shows a selection of curves from the aforementioned comparison. The shorter-wavelength ROI I favors structures with less hydrogen bonding between the nitrogen atoms of trans-azobenzene and the OH group of the ethanol solvent. This is evident from the reduction of hydrogen and oxygen in the curves with an azobenzene N at the origin (Figure 3a,b). This is accompanied by an increase of carbon near the nitrogen center, which we interpret as rotation of the ethanol molecule. Moreover, a blueshift is indicative of shorter N–N and CN=Northo-carbon bond lengths. The simultaneous reduction of N–C curves in short ranges (Figures 3a,b) indicates a shift towards longer N–CN=N bond lengths at the azobenzene C–N=N–C bridge. We note that other notable structural changes are implied as well, and refer to the Supplementary Information for the complete LMBTR vectors. A minor effect may arise from, for example, bending of the N=N–C angles typically associated with the ”inversion pathway” of the photodynamics [15]. To confirm this, we calculated the Pearson correlation coefficient between the ROI values and the angles, finding a slight correlation of around $\sim$0.1–0.2. On the other hand, we note that the C–N=N–C dihedral, associated with the ”rotation pathway” [15], has negligible correlation with the ROI values.
@@
We performed a similar analysis on the ROI value sum in terms of the corresponding latent coordinate $t_{2}$. The results, shown in the Supplementary Information, indicate that hydrogen bonding and the bond lengths have a much lesser effect. Instead, the ROI sum is affected by, for example, twisting of the molecule and an increase in the amount of the ethanol solvent near the very tips of the carbon rings. The result can be affected by a molecule being excluded from the explicitly treated cluster due to the necessitated cutoff distance, but is still indicative of a closer ethanol molecule.
Figure 3: Element-wise change in radial number density ($\Delta$RND) implied by a ten-unit rise of latent coordinate $t_{1}$ associated with ROI value difference from a: the nitrogen atoms, and from b: carbon atoms of the C-N=N-C bridge. We denote regions corresponding to the weakening of the hydrogen bonding and possible rotation of the ethanol molecules with an asterisk (*). Additionally, the triangles ($\triangledown$) denote regions indicating shortening of N=N or CN=N–C bonds, and the square ($\Box$) denotes a region indicating an increase in the length of the N–C bond. The curves shown are an average for the two N atoms or the two CN=N atoms.

### v2

@@ Abstract
We report a statistical simulation of the UV–visible absorption spectrum of trans-azobenzene in ethanol solution. Due to intermolecular interactions, the used explicit solvent environment necessitates accounting for numerous transitions for a spectrum covering the two energetically lowest lines, S1 and S2. Furthermore, the spectrum manifests vast variation as a function of the underlying local structure, in conjunction with previous observations for spectra of liquids in the X-ray regime. We disentangle the complex structure–spectrum relationship using a machine learning-based method known as the emulator-based component analysis. This structural decomposition outperforms commonly used principal component analysis in explained spectral variation and reveals a small subset of latent structural variables responsible for the total spectral variance. Among other structural characteristics, blueshifting of the S2 peak occurs with fewer hydrogen bonds with the ethanol solvent, and a contracted N=N bond within the C–N=N–C bridge. The observed structural dependence of the absorption spectrum thus implies an overrepresentation of certain structural classes after a photoexcitation, potentially significant for the subsequent nuclear dynamics, photophysics, and photochemistry.
I Introduction
A liquid system exhibits a wealth of local atomistic configurations, manifested as significant variation in e.g. its ensemble-averaged X-ray [31, 24, 19, 21, 9] and UV–visible [26, 5, 6] spectra. This underlying diversity is ignored when evaluating molecular properties for a single representative structure, and is probably impossible to capture by a few selected structures and their respective spectra [11]. Statistical variation of spectra implies pre-selection of structures for a spectral response, such as photon absorption at a given wavelength. At the same time, this structure-dependent response allows for structural characterization if the spectrally relevant structural degrees of freedom can be identified. In a liquid system, this is not trivial. To this end, extensive simulations allow an insight into the structure–spectrum relationship when combined with analysis techniques such as the recently introduced emulator-based component analysis (ECA) [20, 28, 7, 8]. This method is used to identify the characteristics of the input $\mathbf{x}$, that are decisive for the response $\mathbf{y}(\mathbf{x})$, thus simplifying the inverse problem posed by structural ($\mathbf{x}$) interpretation of spectra ($\mathbf{y}$). Such target-variance-driven dimensionality reduction is also useful when the outcome $\mathbf{y}$ depends predominantly on a small, unknown subset of features in a high-dimensional $\mathbf{x}$. The ECA method requires approximating $\mathbf{y}(\mathbf{x})$ by a reasonably accurate and computationally efficient emulator $\mathbf{y}_{\mathrm{emu}}(\mathbf{x})$, for which a neural network is a contemporary choice. Potential applications of the protocol include, but are not limited to, statistical phenomena in molecular liquids and in their spectra, the topic of this work.
In this work, we study the UV–visible spectrum of trans-azobenzene in ethanol solution by combining molecular dynamics (MD) simulations with subsequent spectral calculations and dimensionality reduction procedures. We find that the dominant structural variation in this complex system, derived by principal component analysis (PCA), does not explain the intensity behavior within key spectral regions. Instead, decomposition by ECA captures the spectral variance with just a few latent structural degrees of freedom, that cover a small fraction of the total structural spread. With this low-dimensional representation, we identify the most significant structural characteristics behind a shift in the S2 peak. Finally, we discuss the implications of these findings for processes occurring after photoexcitation.
II Methods
We sampled 30 000 structures from extended tight binding MD (xTBMD) using the GFN1-xTB method [12, 2] implemented in the CP2K software [16] v. 2024.2. In the simulations, a trans-azobenzene molecule was dissolved in 159 ethanol molecules in a cubic box of $L$ = 25.04 Å for the density 0.7840 g/cm3 interpolated for pure ethanol liquid at 300 K [15]. After equilibration of 20 ps, we sampled the canonical ensemble at 300 K with a 10 fs interval from a production run of 300 ps (for validation, see SI). We used a global Nosé–Hoover thermostat [23, 22] with chain length 3, a time constant $\tau$ of 1000 fs, and an MD timestep of 0.5 fs.
@@ II Methods
For evaluation of spectra consisting of 50 transitions on each of the 30 000 structures, we applied time-dependent density functional theory (TD-DFT) implemented in ORCA v.6.0.0 [18]. We explicitly included ethanol within the cutoff radius of 3.0 Å, and simulated this cluster in a conductor-like polarizable continuum model [10]. We used B3LYP exchange-correlation potential [17, 3], def2-SVP basis set [29] and def2/J auxiliary basis set [30] in the spectral simulations. We evaluated the cross section $\sigma_{i}$ of each transition $i$ for a randomly oriented sample using the transition matrix element in the velocity form. We then convolved this stick spectrum to differential cross section $\mathrm{d}\sigma/\mathrm{d}E$ in terms of energy $E$ by a Gaussian function (full-width at half maximum 0.414 eV), together with a constant shift of 0.300 eV in energy. We obtained these required parameters from a fit of the the computational ensemble-averaged spectrum to an experiment by Nägele and coworkers [Nägele1997]. Finally, we transformed the differential cross section into wavelength $\lambda$ scale for $\mathrm{d}\sigma/{\mathrm{d}\lambda}$. We use a constant convolution width in energy, because a constant width in wavelength makes the lines artificially broad for short wavelengths.
@@
To capture the structure–spectrum relationship in these data, we train a neural network emulator and use it to iteratively find the ECA vectors for identifying spectrally decisive structural characteristics in the system. As the raw atomic coordinates from simulations are suboptimal for machine learning, we preprocess the structural information using a local version of the many-body tensor representation (LMBTR) [13]. The LMBTR descriptor consists of Gaussian-smeared element-wise radial distance distributions for multiple centers in the system. As these centers, we use all the nitrogen and carbon sites in the azobenzene molecule, and eight additional virtual centers located perpendicular to a plane defined by nearby atoms in the azobenzene molecule. The first four virtual centers lie $\pm 2$ Å from each nitrogen atom. The latter four centers lie $\pm 2$ Å perpendicular from the arithmetic mean position of the para and ortho carbon atoms of each ring. For each structure, the encoding procedure resulted in a vector $\mathbf{d}$, which we z-score standardized to produce $\mathbf{\tilde{d}}$ used in machine learning and ECA. To establish a well-performing LMBTR–neural network combination, we ran a joint search [7] including both LMBTR and neural network hyperparameters, and chose the best-performing combination. The final model had an R2 score (coefficient of determination) of 0.945 on a test set of 5 000 points, split from the data pool of 30 000 points prior to model selection, and thus previously unseen by the model. For more details, we refer to SI.
@@
Figure 1: Results of the spectrum calculations and variance captured by different dimensionality reduction techniques. a: Simulated ensemble average UV–visible absorption spectrum for trans-azobenzene in ethanol, together with an experiment reproduced from Ref. Nägele1997. The blue-shaded area depicts the standard deviation of the data set, indicating significant statistical variation. The gray-shaded areas depict the two regions of interest (ROIs I and II) defined for the subsequent analysis. Computational results have been shifted to match the experiment. Additionally, the figure presents an example local structure prepared with Jmol [14]. b: Spectral R2 score as a function of the rank of emulator-based component analysis (ECA) and principal component analysis (PCA), evaluated using the neural network emulator after the respective dimensionality reduction. The first two ECA components explain a large portion of the spectral ROI variance, while the higher components show rapidly diminishing R2 score gains. Structural decomposition by PCA covers significantly less spectral variance for the same rank. c: Structural R2 score as a function of decomposition rank. The ECA vectors cover a small fraction of the structural variance compared to structural PCA. The result indicates that explaining structural variation alone does not explain the spectral variation.
III Results
Figure 1a shows the simulated ensemble-averaged spectrum together with the experimental one by Nägele and coworkers [Nägele1997]. We apply the molecular notation S1 and S2 for the characteristic features in this spectrum, but note that in our simulation for the gas phase in the optimized geometry, the S2 line already has three underlying states. In liquid, we find the photoexcited states of trans-azobenzene to be mixed with states of the ethanol, which results in a wealth of possible final states to be excited – and to be simulated with explicit solvation. In our data, the average number of transitions within lines S1 (390 nm–520 nm) and S2 (270 nm–390 nm) were 0.9 and 14.1, respectively. The S1 feature is almost unobservable, and we focus on the S2 peak. We observe significant statistical variation between the spectra of individual local structures, as shown by the standard deviation $\sigma$ in Figure 1a. For the following analysis, we define two regions of interest (ROI), split at the maximum of the S2 peak. Then, we study the mean differential cross section within each ROI as a function of the structure. This practice makes the analysis focus on overall features of the contributing spectra within the peak, instead of their fine details, that are more prone to error due to the used approximations. We first train a neural network to predict the two z-score standardized ROI values for a structure, and then utilize this emulator in in the ECA procedure for dimensionality reduction to identify the decisive structural characteristics behind the observed spectral variation.
@@ III Results
The ECA method is analogous to PCA with the key difference being that the optimization target of the former is the explained ROI intensity variance. As seen in Figure 1b, rapid convergence towards the emulator performance limit is observed when emulation is carried out on data points projected onto the low-dimensional subspace. In contrast, applying PCA to the structural data points $\mathbf{\tilde{d}}$ aims at finding the decisive components of structural variance, not necessarily reflected by the spectral response as seen in Figure 1b. Vectors from structural PCA cover significantly less spectral variance after the corresponding evaluation. Perhaps even more notably, the dominant structural degree of freedom is spectrally irrelevant. The different optimization goal of the methods is also evident in the structural R2 score as a function of the decomposition rank, presented in Figure 1c, which shows PCA to surpass the ECA decomposition. Thus, the majority of the spectral ROI intensity variance is explained by a minority of structural variance. We note that another method, partial least squares fitting, has previously been compared to ECA, with the latter excelling in this comparison [20].
@@
Next, we turn to the analysis of the first two ECA component vectors, that indicate the structural changes responsible for most spectral ROI intensity variation. We first analyze the spectral effect, and then the associated structural changes. We define two spectral properties: (i) ROI value difference (ROI I – ROI II), and (ii) total ROI value sum. These properties can be directly linked in the two latent coordinates $t_{1}$ and $t_{2}$, as indicated in Figure 2. The spectrally most significant structural degree of freedom $t_{1}$ corresponds to the ROI value difference (Figure 2a), whereas the second most significant degree of freedom $t_{2}$ captures the ROI value sum (Figure 2d). The two properties are mostly independent of each other, illustrated by the low cross-correlation in Figures 2b and 2c. From Figures 2a and 2d we conclude that ROI I intensity is high when $t_{1}$ and $t_{2}$ values are high. Similarly, ROI II intensity is high when $t_{1}$ value is low and $t_{2}$ high. Thus, the structural changes causing a blueshift of the S2 line are indicated by the latent coordinate $t_{1}$, as the effects from the second ECA component cancel out. For a structural interpretation of such effect, we calculate the difference 10 $\mathbf{v}_{1}$ for a well-expected change in this coordinate, and then transform the respective ECA component vector to the original descriptor space. A further division by the squared distance to a corresponding center yields site-wise and element-wise curves with radial behavior proportional to that of the radial distribution function.
@@
Figure 2: The relationship between spectral properties and the latent coordinates from ECA using the test set with the corresponding Pearson’s correlation coefficients $r$. a: The difference between the two ROI values is highly correlated with the spectrally most significant structural variable $t_{1}$. b: The correlation between ROI value difference and the second most significant structural variable $t_{2}$ is negligible. c: The total ROI value sum is weakly correlated with $t_{1}$, but d: highly correlated with $t_{2}$. For details, see text.
@@
Figure 3 shows a selection of curves from the aforementioned comparison. The shorter-wavelength ROI I favors structures with less hydrogen bonding between the nitrogen atoms of trans-azobenzene and the OH group of the ethanol solvent. This is evident from the reduction of hydrogen and oxygen in the curves with an azobenzene N at the origin. This is accompanied by an increase of carbon near the nitrogen center, which we interpret as rotation of the ethanol molecule. Moreover, a blueshift is indicative of shorter N=N and CN=Northo-carbon bond lengths. The simultaneous reduction of N–C curves in short ranges indicates a shift towards longer N–CN=N bond lengths at the azobenzene C–N=N–C bridge. The results were consistent across ten of the best-performing descriptor–neural network architecture combinations.
While the ECA procedure is carried out without a priori assumptions, the sanity of the obtained results can be confirmed a posteriori by evaluating the mean spectra for structures differing in the according characteristics. Using the whole data set, Figure 4 shows a blueshift of the S2 peak for structures with shorter N=N bond lengths and for structures with less hydrogen bonds for N atoms, evaluated using geometric criteria [21]. We note that other notable structural changes are implied as well, and refer to the SI for the complete list.
[…]
Figure 4: Validation of the ECA results evaluated over the whole data set. a: The mean spectra of structures with N=N bond length above and below the average value. b: The mean spectra for structures with a given number of hydrogen bonds accepted by the azo group, calculated with criteria described in Ref. [21]. The inspection is motivated by ECA first finding these characteristics.
@@
Additionally, we performed a corresponding analysis for the total ROI value sum, the results of which are presented in the SI. This spectral property appears to be affected, for example, increase in the amount of the ethanol solvent near the very tips of the carbon rings. The result can be affected by a molecule being excluded from the explicitly treated cluster due to the necessitated cutoff distance, but is still indicative of a closer ethanol molecule. The analysis also suggests that numerous other less interpretable structural features affect the ROI value sum.
We repeated the calculations using the PBE exchange–correlation potential [25], which in this case required evaluating 100 transitions. The aforementioned structural conclusions remain unaffected in the subsequent analysis, performed in a manner analogous to the presented one and reported in SI.
IV Discussion